# Matching Free Depth Recovery from Structured Light

### v1

@@ 3.3 Loss Functions
In our scenario, the volumetric density along each sampled light ray should be singular peaked, as each ray from camera should intersect only once with the object’s surface. Besides, due to the lack of camera views in our task, there may be multiple peaks along the projecting ray to produce one precise color, resulting in lots of floaters in the reconstructed geometry. Thus we apply a distortion loss proposed by Mip-NeRF 360 [3]. For a ray with $K$ sampled points, this loss is defined as
@@
where $(s_{i+1}-s_{i})$ is the length, $(s_{i}+s_{i+1})/2$ is the midpoint of the $i$-th query interval and $w_{i}=T_{i}\alpha_{i}$. The first term minimizes the weighted distances between all pairs of interval midpoints, and the second term minimizes the weighted size of each individual interval. As a result, the trained geometry is forced to fit our task. The total distortion loss is
@@
The surface color loss is similar to photometric constraints which are applied in matching-based techniques [10, 50], where colors are warped between images using pixel depth information. It can enhances geometry quality and overall performance, especially in few-shot scenarios [6, 24, 43]. Our approach prioritizes geometric performance over photorealism. Therefore, enforcing density constraints helps the network generate more accurate 3D shapes without introducing rendering ambiguities. We provide a schema to illustrate the difference between rendered color loss and surface color loss in Fig. 3.
@@
Once the training process is finished, we extract the full depth map $\mathbf{D}$ of input scene using Eq.(11).
3.4 NDC Parameterization
Our task focuses on front-facing scenes, as there’s only one camera in a monocular structured light system. Inspired by the original NeRF [22], we define our voxel grid in normalized device coordinates(NDC) space. By warping an infinitely deep camera frustum into a bounded cube, where distance along the z-axis corresponds to disparity (inverse distance), NDC efficiently reallocates the voxel grid’s density in a way that aligns with the geometry of perspective projection.
@@ 3.4 NDC Parameterization
4 Experiments
4.1 Datasets
We use patterns as color constraints for training the voxel grid to represent 3D scene geometry, eliminating the need for explicit matching techniques typically applied to pattern sets. Thus, we simply use a set of randomly generated, array-based patterns. The projector space is divided into uniform squares, with each square randomly assigned either black or white. To generate our pattern sets, we use unit squares of varying sizes. We select pattern lengths of 20, 10, and 5 pixels to generate a total of 6 patterns (two patterns per scale) for experiments, same as in SL-SDF [27].
To evaluate our approach and compare it with existing methods, we first assess the accuracy of depth estimation on synthetic scenes. These scenes are rendered using the tool provided by CTD [28], following the same experimental setup as CTD and using objects from the ShapeNet dataset [5]. A total of 50 different scenes were rendered to create our synthetic dataset. The camera’s intrinsic parameters and the projector’s settings are similar to those used in SL-SDF [27] to facilitate later comparison. Additionally, we test our method on a real dataset provided by [27], which includes 6 scenes with estimated ground truth depths.
N-PMP(6)
H-PMP(6)
CGC(6)
CGC(7)
GC(8)
GC(9)
SL-SDF(6)
Ours(6)
Figure 4: Pattern sets used by different methods. Note that they are vertically repetitive.
Synthetic
Real
GT
N-PMP(6)
H-PMP(6)
CGC(6)
CGC(7)
GC(8)
GC(9)
SL-SDF(6)
Ours(6)
Figure 5: Visualization of estimated depths and error maps by different methods on both synthetic and real scene. The disparity differences between estimated results and ground truth are displayed in the second and fourth rows.
@@ 4.1 Datasets
err:avg $O(0.1)$ $O(0.5)$ $O(1)$ $O(2)$ err:avg $O(0.1)$ $O(0.5)$ $O(1)$ $O(2)$
N-PMP(6) [51] 87.461 1.80 1.62 1.32 0.65 513.855 68.99 67.13 62.22 43.01
H-PMP(6) [39] 40.192 0.49 0.15 0.15 0.14 56.635 14.52 3.11 2.12 0.40
CGC(6) [41] 83.594 1.32 0.90 0.84 0.75 46.857 7.05 5.29 0.81 0.50
CGC(7) [41] 69.538 1.02 0.89 0.49 0.38 38.849 6.00 3.90 1.52 0.53
GC(8) [1, 40] 56.567 0.65 0.61 0.56 0.45 18.941 4.92 3.17 2.20 0.63
GC(9) [1, 40] 43.715 0.55 0.50 0.41 0.23 16.171 4.30 2.19 1.23 0.56
SL-SDF(6) [27] 42.056 0.49 0.39 0.39 0.39 9.376 3.14 0.82 0.45 0.31
Ours 13.767 0.08 0.06 0.06 0.06 9.153 2.87 0.92 0.47 0.40
Table 1: The experiment results compared with classic methods on both synthetic and real scenes. The unit of length for "err:avg" here is millimeter. Results on the synthetic dataset represent the average values across 50 scenes, and those on the real dataset are an average of 6 scenes. We color code the best and second best.
@@
err:avg $O(0.1)$O(0.5) \$O(1) \$O(2) err:avg $O(0.1)$O(0.5) \$O(1) \$O(2)
$\mathcal{L}_{\text{photo}}$ 35.844 0.21 0.12 0.11 0.11 38.712 7.13 4.69 1.91 0.54
$\mathcal{L}_{\text{photo}}+\mathcal{L}_{\text{dist-total}}$ 27.210 0.15 0.12 0.11 0.11 33.948 6.74 3.81 1.49 0.52
$\mathcal{L}_{\text{photo}}+\mathcal{L}_{\text{surface}}$ 18.573 0.10 0.07 0.06 0.06 14.394 3.38 1.51 0.98 0.65
$\mathcal{L}_{\text{photo}}+\mathcal{L}_{\text{dist-total}}+\mathcal{L}_{\text{surface}}$ 13.767 0.08 0.06 0.06 0.06 9.153 2.87 0.92 0.47 0.40
Table 2: Ablation studies conducted to assess the impact of the rendered color loss, the distortion loss and the surface color loss. Numbers here also represent the average results of all scenes. We color code the best and second best.
4.2 Implementation Details
We use consistent hyperparameters across all scenes. We divide the xy-plane of the NDC cube to $256\times 256$ and discretize the length along the z-axis into $256$ different disparity values, resulting in an expected voxel count of $M=256^{3}$. We set $\alpha^{(\text{init })}=10^{-2}$ and the point sampling step size is chosen as half of the voxel size. For training the voxel grid, we use $\lambda_{d}=0.01$ with $\lambda_{s}=0$ for the first 3000 iterations, during which we sample 8192 rays per iteration. The exclusion of surface color loss during this phase is a strategic choice to avoid potential issues with local minima. After the initial phase, we set $\lambda_{s}=1$ and continue training for an additional 29,000 iterations. The entire training process takes approximately 5 minutes on a single NVIDIA GTX 4060.
4.3 Comparisons
We first evaluate our method against four classic decoding-based structured light techniques, as the core principles of temporal-encoding structured light methods have remained unchanged over the past decade. Specifically, we compare with Numerical Phase Measurement Profilometry (N-PMP) [51], Hierarchical Phase Measurement Profilometry (H-PMP) [39], Binary Gray Code (GC) with interpolation between fringes [1, 25, 40], and Complementary Gray Code (CGC) [41]. N-PMP and H-PMP both require six patterns for decoding, and CGC require six or seven patterns. GC method can take different number of patterns for calculation, here we choose eight and nine for comparison. We also compare our method with SL-SDF [27], which is also a matching free approach.  [27] builds a neural signed distance field to represent 3D geometry, and applys a NeuS-based [38] differentiable rendering scheme during training phase. The depth map is then generated through marching cubes algorithm [20] and re-projection from the trained SDF. The specific patterns used by each method are illustrated in Fig.4, with the corresponding pattern counts indicated in parentheses. To make competitive experiments against SL-SDF, our 2D pattern sets are set to the same as theirs. We use both the average L1 loss and the outlier metric to evaluate the recovered depth maps of each method. The definition of percentage of outliers $o(t)$ is from CTD [28], which is the percentage of pixels where the difference between the estimated and ground-truth disparities (inverse depths) is greater than a certain threshold $t$. We summarize our experimental results and show them in Table 1. Results shown in the table represent the average depth error and percentage of outliers on each scene in the dataset. Additionally, we visualize estimated images in Fig. 5. The first row shows visualizations of the disparity maps generated by various methods on our synthetic scenes, while the second row presents the error maps between the estimated disparities and the ground truth. The subsequent two rows display the corresponding results on real data. Here disparities are used for better visualization.
The N-PMP, H-PMP, and CGC methods utilize phase-shifting encoding, which encodes the coordinates into a set of sinusoidal patterns. However, these patterns are periodic, introducing phase ambiguities during decoding. Typically, this issue is addressed by projecting additional patterns with different frequencies. While phase-shifting techniques can achieve high precision with a large number of patterns, they encounter challenges when the pattern count is limited. N-PMP uses two sets of phase patterns with shorter wavelengths, making it sensitive to errors in the decoded phase values. In contrast, H-PMP adds an extra set of phase patterns with a wavelength equivalent to the image width, but it remains susceptible to shading effects. CGC utilizes binary gray-code patterns, but the long wavelengths of these patterns reduce its decoding accuracy. GC relies entirely on binary gray-code patterns, which can lead to inaccuracies between the projected binary fringes
In contrast to traditional methods, both SL-SDF [27] and our approach utilize 2D binary patterns without the need for pattern matching. Experiments demonstrate that both methods are capable of generating smooth surfaces. However, SL-SDF struggles on certain synthetic datasets, particularly those containing sharp surface edges or complex geometries. This limitation stems from its use of signed distance fields (SDFs). While SDFs effectively enforce surface smoothness, they face difficulties in modeling sharp geometric features, as shown in Fig. 6. This issue becomes more pronounced in structured light depth reconstruction scenarios, where textures and color information are reduced. As a result, fine details and sharp edges may be smoothed out or inaccurately reconstructed, impacting the overall fidelity of the recovered depth. In contrast, our voxel-grid-based representation offers significant advantages in handling such scenarios. Unlike SDFs, voxel grids explicitly model scene properties at discrete volumetric locations, enabling more accurate encoding of abrupt geometric changes and providing greater flexibility in capturing sharp edges. The surface color loss and distortion loss we apply ensure the smoothness of our estimated depths while maintaining geometric accuracy.
SL-SDF(6)
Ours(6)
Figure 6: Difference of the estimated depth map at object edges between our method and SL-SDF [27]. SDFs tend to generate smooth surfaces, making it challenging to handle sharp edges.
4.4 Ablation Studies
To evaluate the impact of the loss terms introduced in Section 3.3, we conduct experiments with different combinations of loss functions on our synthetic datasets. The results, presented in Table 2, demonstrate that the primary photometric loss, while effective in reconstructing the overall scene geometry, struggles to capture fine-grained details accurately. The addition of surface color loss significantly improves the representation of these finer geometric features. Furthermore, the distortion loss, which enforces a singular peak along each sampled ray, further enhances the final depth estimation performance.
Figure 7: Comparison with SL-SDF on training time and accuracy. We use the same set of 6 patterns as in their approach. Our training converges after 5 minutes while theirs need over 14 minutes.
In order to demonstrate the advantages of our voxel grid training schedule for structured-light-based depth recovery, we make further comparison with SL-SDF [27], which uses signed distance fields (SDFs) for a similar procedure. We explore the relationship between training time and depth estimation accuracy for both methods. As shown in Fig. 7, our method requires significantly less training time to achieve the same level of depth reconstruction accuracy. This efficiency can be attributed to several factors, with one of the main reasons being the inherent differences in iteration speed between explicit and implicit scene representations. Specifically, with the same ray sampling batch size, each iteration in our method is approximately 20 times faster than in SL-SDF.
Furthermore, we examine the effect of the number of projected patterns on the performance of our method. We first increase the pattern count to 9(with 3 patterns per scale) to demonstrate the full result, and then remove one of them after each test. The results demonstrate that our voxel-based training process achieves satisfactory performance with as few as six patterns, with no significant improvement observed by further increasing the number of patterns. In addition, our method produces promising results when using only four patterns.
GT
Figure 8: Visualized results of varying pattern numbers. The number below each picture stands for the amount of patterns we used to capture scenes and train our voxel grid. The last one shows the ground truth depth map.
Pat 2 3 4 5 6 7 8
Err 203.9 31.7 18.3 16.9 15.3 15.2 15.0
Table 3: Ablation study on different pattern numbers within our framework. Errors stand for the average depth error across all test scenes. We verify 10 synthetic scenes for this experiment.
5 Conclusion

### v2

@@ III-C Loss Functions
Figure 4: Visualization of our photometric loss $\mathcal{L}_{\text{photo }}$, distortion loss $\mathcal{L}_{\text{dist }}$, and surface color loss $\mathcal{L}_{\text{surface}}$. The photometric loss encourages the rendered colors to match the captured image, ensuring overall appearance consistency. In contrast, both the distortion loss and the surface color loss promote a single dominant peak in the volume density distribution along each ray, thereby improving surface localization and reducing ambiguity in depth estimation.
In our scenario, the volumetric density along each sampled light ray should be singular peaked, as each ray from the camera should intersect only once with the object’s surface. Besides, due to the lack of camera views in our task, there may be multiple peaks along the projecting ray to produce one precise color, resulting in lots of floaters in the reconstructed geometry. Thus, we apply a distortion loss proposed by Mip-NeRF 360 [48]. For a ray with $K$ sampled points, this loss is defined as
@@
where $(s_{i+1}-s_{i})$ is the length, $(s_{i}+s_{i+1})/2$ is the midpoint of the $i$-th query interval and $w_{i}=T_{i}\alpha_{i}$. The first term minimizes the weighted distances between all pairs of interval midpoints, and the second term minimizes the weighted size of each individual interval. As a result, we force the trained geometry to fit our task. The total distortion loss is
@@
The surface color loss is similar to photometric constraints, which are applied in matching-based techniques [49, 42], where colors are warped between images using pixel depth information. It can enhance geometry quality and overall performance, especially in few-shot scenarios [50, 22, 51]. Our approach prioritizes geometric performance over photorealism. Therefore, enforcing density constraints helps the voxel grid generate more accurate 3D shapes without introducing rendering ambiguities. We provide a schema to illustrate the difference between our losses in Fig. 4.
@@
Once the training process is finished, we extract the full depth map $\mathbf{D}$ of the input scene using Eq.(11).
III-D NDC Parameterization
Our task focuses on front-facing scenes, as there’s only one camera in a monocular structured light system. Inspired by the original NeRF [19], we define our voxel grid in normalized device coordinates(NDC) space. The transformation from the camera frustum to the NDC space in our process is illustrated in Fig. 5. The standard 3D perspective projection matrix for homogeneous coordinates is given by:
$M=\left(\begin{array}[]{cccc}\frac{n}{r}&0&0&0\\ 0&\frac{n}{t}&0&0\\ 0&0&\frac{-(f+n)}{f-n}&\frac{-2fn}{f-n}\\ 0&0&-1&0\end{array}\right)$ (15)
where $n$ and $f$ represent the near and far clipping planes, and $r$ and $t$ are the right and top bounds of the frustum at the near clipping plane. Given a homogeneous point $(x,y,z,1)^{T}$, we apply the transformation by left-multiplying the point by the matrix $M$ and then dividing out the fourth coordinate to get the projected point
$(-\frac{nx}{rz},-\frac{ny}{tz},\frac{(f+n)}{f-n}+\frac{2fn}{(f-n)z})^{T}$ (16)
The projected point is now in normalized device coordinate (NDC) space, where the original viewing frustum is mapped to the cube $[-1,1]^{3}$. By mapping the top-right pixel on the image plane to the top-right corner of the near plane, we obtain:
$\frac{f_{x}}{c_{x}}\frac{r}{n}=1,\quad\quad\frac{f_{y}}{c_{y}}\frac{t}{n}=1$ (17)
Where $n$ represents the chosen near clipping plane, $f_{x},f_{y}$ and $c_{x},c_{y}$ are focal lengths and principal points of the camera. Therefore, $r$ and $t$ can be computed using the camera parameters:
$r=\frac{nc_{x}}{f_{x}},\quad\quad t=\frac{nc_{y}}{f_{y}}$ (18)
We further set $f\to\infty$ in our approach, deriving the relationship between world coordinate $(x,y,z)^{T}$ and its corresponding NDC $(x_{*},y_{*},z_{*})^{T}$
$\begin{split}&x=2n\frac{c_{x}}{f_{x}}\frac{x_{*}}{1-z_{*}}\\ &y=2n\frac{c_{y}}{f_{y}}\frac{y_{*}}{1-z_{*}}\\ &z=\frac{2n}{z_{*}-1}\end{split}$ (19)
Here $z\in(-\infty,-n]$ and $z_{*}\in[-1,1)$ as the camera is looking in the $-z$ direction.
Figure 5: Transformation from camera frustum in world space to NDC space. (a) world coordinates. (b) normalized device coordinates. We define the near and far planes as the depth boundaries for this transformation.
By warping an infinitely deep camera frustum into a bounded cube, where distance along the z-axis corresponds to disparity (inverse distance), NDC efficiently reallocates the voxel grid’s density in a way that aligns with the geometry of perspective projection.
@@ III-D NDC Parameterization
IV Experiments
IV-A Datasets
TABLE I: Camera and projector parameters used in real-world scenes
Device $f_{x}$ $f_{y}$ $c_{x}$ $c_{y}$ Baseline
Camera 1181.76 1179.92 639.50 511.50 209.39
Projector 2013.30 2016.43 699.16 755.26
We use patterns as color constraints for training the voxel grid to represent 3D scene geometry, eliminating the need for explicit matching techniques that are applied to pattern sets and captured images. Thus, we simply use a set of randomly generated 2D binary patterns. The projector space is divided into uniform squares, with each square randomly assigned either black or white. To generate our pattern sets, we use unit squares of varying sizes. We select pattern lengths of 20, 10, and 5 pixels to generate a total of 6 patterns (two patterns per scale) for experiments.
To evaluate our approach and compare it with existing methods, we first assess the accuracy of depth estimation on synthetic scenes. These scenes are rendered using the tool provided by CTD [13], following the same experimental setup as CTD and using objects from the ShapeNet dataset [52]. A total of 50 different scenes are rendered to create our synthetic dataset. Additionally, we test our method on real scenes provided by SL-SDF [27], which includes six scenes with estimated ground truth depths. Device parameters of these scenes are shown in Table I.
TABLE II: Pattern Types, Number of Patterns, and Pretraining Requirements of Different Methods
N-PMP H-PMP CGC GC CTD GigaDepth SL-SDF Ours
Binary Patterns(1D) - - 4 9 - - - -
Sinusoidal Patterns(1D) 6 6 3 - - - - -
Binary Patterns(2D) - - - - - - 6 6
Speckle Patterns(2D) - - - - 1 1 - -
Pretraining Datasets - - - - required required - -
IV-B Implementation Details
We use consistent hyperparameters across all scenes. We divide the xy-plane of the NDC cube into $256\times 256$ and discretize the length along the z-axis into $256$ different disparity values, resulting in an expected voxel count of $M=256^{3}$. We set $\alpha^{(\text{init })}=10^{-2}$ and the point sampling step size is chosen as half of the voxel size. For training the voxel grid, we use $\lambda_{d}=0.01$ with $\lambda_{s}=0$ for the first 3000 iterations, during which we sample 8192 rays per iteration. The exclusion of surface color loss during this phase is a strategic choice to avoid potential issues with local minima. After the initial phase, we set $\lambda_{s}=1$ and continue training for an additional 29,000 iterations. The entire training process takes approximately 5 minutes on a single NVIDIA GTX 4060.
IV-C Comparisons
Figure 6: Visualization of estimated depth maps and corresponding error maps from different methods on synthetic scenes.
Figure 7: Visualization of estimated depth maps and corresponding error maps from different methods on real scenes.
TABLE III: Quantitative Comparison of Depth Recovery Performance Across Synthetic and Real-World Scenes
@@ IV-C Comparisons
MAE(mm) $O(0.1)$ $O(0.5)$ $O(1)$ MAE(mm) $O(0.1)$ $O(0.5)$ $O(1)$
N-PMP(6) [37] 29.431 2.05 1.85 1.71 513.855 68.99 67.13 62.22
H-PMP(6) [36] 30.787 2.16 1.09 0.61 56.635 14.52 3.11 2.12
CGC(7) [53] 33.445 2.04 1.72 1.59 38.849 6.00 3.90 1.52
GC(8) [32, 33] 22.539 1.36 1.03 0.96 18.941 4.92 3.17 2.20
GC(9) [32, 33] 19.087 1.02 0.62 0.57 16.171 4.30 2.19 1.23
CTD(1) [13] 15.613 0.72 0.10 0.10 296.560 18.00 16.24 16.01
GigaDepth [15] 14.176 0.10 0.06 0.06 18.629 2.98 1.47 0.73
SL-SDF(6) [27] 90.070 5.51 2.14 2.14 9.376 3.14 0.82 0.45
Ours 13.767 0.08 0.06 0.06 9.153 2.87 0.92 0.47
We first evaluate our method against four classic decoding-based structured light techniques. Specifically, we compare with Numerical Phase Measurement Profilometry (N-PMP) [37], Hierarchical Phase Measurement Profilometry (H-PMP) [36], Binary Gray Code (GC) with interpolation between fringes [32, 33], and Complementary Gray Code (CGC) [53]. N-PMP and H-PMP require six projected patterns, and CGC requires seven. The GC method can take a different number of patterns for calculation, here we choose eight and nine for comparison. We then compare our method with two learning-based methods, Connecting the Dots(CTD) [13] and GigaDepth [15]. CTD formulates depth estimation as a convolutional neural network task, while GigaDepth models it as a regression problem and further decomposes the regression into smaller classification sub-tasks using multi-layer perceptrons (MLPs). Both methods require pretraining on hundreds of synthetically generated scenes to achieve satisfactory performance. Besides, We also compare our method with SL-SDF [27], which is a matching-free approach.  [27] builds a neural signed distance field to represent 3D geometry, and applys a NeuS-based [23] differentiable rendering scheme during training phase. The depth map is then generated through the marching cubes algorithm [29] and re-projection from the trained SDF.
We illustrate the types and numbers of patterns used by each method with the requirements of pretraining on generated datasets in Table II. To ensure a fair comparison with SL-SDF, we use the same set of 2D projected binary patterns as employed in their approach. We use both the mean absolute error and the outlier metric to evaluate the recovered depth maps of each method. When calculating MAE for classic decoding-based methods, we set the depth value of outliers to the mean value of the estimated depth map to minimize their influence. The definition of the percentage of outliers $o(t)$ is from CTD [13], which is the percentage of pixels where the difference between the estimated and ground-truth disparities (inverse depths) is greater than a certain threshold $t$. We summarize our experimental results and show them in Table III. Results shown in the table represent the average depth error and percentage of outliers across all scenes in the dataset(50 synthetic scenes and six real scenes). Additionally, we visualize the recovered depth maps and error maps by different methods in Fig. 6 and Fig. 7. GT stands for the ground truth. Here, we use disparities for better visualization. We also show edge details from different methods in Fig. 9 for further comparisons.
The N-PMP, H-PMP, and CGC methods employ phase-shifting encoding, which represents pixel coordinates using sets of sinusoidal patterns. However, due to their periodic nature, these patterns introduce phase ambiguities during decoding. This issue is typically mitigated by projecting additional patterns at varying spatial frequencies. While phase-shifting techniques can achieve high accuracy with a sufficient number of patterns, they become less reliable when the number of patterns is limited, especially in captured real scenes.
N-PMP utilizes two sets of phase patterns with shorter wavelengths, which increases its sensitivity to phase decoding errors. H-PMP supplements traditional phase patterns with an additional pattern set having a wavelength equal to the image width, but remains vulnerable to shading effects and intensity variation. CGC incorporates binary Gray-code patterns to resolve ambiguities; however, the long wavelengths of these binary patterns limit their decoding precision. GC, which relies solely on binary Gray-code patterns, is prone to errors between adjacent binary fringes due to limited spatial resolution.
CTD and GigaDepth perform well on synthetic datasets, benefiting from carefully designed training scenes with ideal lighting and noise-free conditions. However, their effectiveness diminishes when applied to real captured scenes, primarily due to a significant domain gap between synthetic and real-world data. Moreover, both methods struggle to preserve fine geometric details at object boundaries. CTD’s reliance on convolutional architectures may limit its ability to handle high-frequency variations, while GigaDepth’s discretization of the regression task can lead to quantization artifacts near depth discontinuities. As a result, their depth estimations at object edges often appear overly smoothed or inaccurate, which is particularly problematic in applications requiring precise surface reconstruction.
Both SL-SDF [27] and our approach employ 2D binary patterns and eliminate the need for explicit pattern matching. Experimental results indicate that both methods can generate smooth surface reconstructions. However, SL-SDF exhibits degraded performance in certain synthetic scenes, particularly those containing sharp edges or intricate geometric structures. This limitation arises from its use of signed distance fields (SDFs) as the underlying geometry representation.
While SDFs inherently promote surface smoothness due to their continuous and differentiable formulation, they are less effective at preserving fine-scale geometric discontinuities. This drawback is especially evident in structured light scenarios, where high-frequency textures and rich color cues are largely absent. Consequently, sharp features such as object boundaries may be smoothed out or inaccurately reconstructed, as illustrated in Fig. 9, ultimately reducing the geometric fidelity of the recovered depth map.
In contrast, our method leverages an explicit voxel grid representation, which models scene properties at discrete volumetric locations. This design offers enhanced flexibility in encoding abrupt changes in geometry and enables more accurate recovery of sharp surface transitions. Furthermore, the incorporation of surface color loss and distortion loss in our training framework enforces surface consistency while preserving local geometric details, achieving a favorable balance between smoothness and accuracy.
Figure 8: We adopt the same set of six projected patterns as used in SL-SDF for a fair comparison. Our method achieves convergence within 5 minutes of training, whereas SL-SDF requires more than 14 minutes. Despite the significantly reduced training time, our method attains comparable or superior accuracy.
In order to demonstrate the advantages of our voxel grid training schedule for structured-light-based depth recovery, we make further comparisons with SL-SDF [27], which uses signed distance fields (SDFs) for a similar procedure. We explore the relationship between training time and depth estimation accuracy for both methods. As shown in Fig. 8, our method requires significantly less training time to achieve the same level of accuracy of recovered depths. This efficiency can be attributed to several factors, with one of the main reasons being the inherent differences in iteration speed between explicit and implicit scene representations. Specifically, with the same ray sampling batch size, each iteration in our method is approximately 20 times faster than in SL-SDF.
Figure 9: Visualization of the differences in estimated depth maps at object boundaries between our method and prior approaches.
TABLE IV: Ablation studies to analyze the impact of the rendered color loss, distortion loss, and surface color loss on the accuracy of depth recovery
@@
MAE(mm) $O(0.1)$ $O(0.5)$ $O(1)$ MAE(mm) $O(0.1)$ $O(0.5)$ $O(1)$
$\mathcal{L}_{\text{p}}$ 35.844 0.21 0.12 0.11 38.712 7.13 4.69 1.91
$\mathcal{L}_{\text{p}}+\mathcal{L}_{\text{d-t}}$ 27.210 0.15 0.12 0.11 33.948 6.74 3.81 1.49
$\mathcal{L}_{\text{p}}+\mathcal{L}_{\text{s}}$ 18.573 0.10 0.07 0.06 14.394 3.38 1.51 0.98
$\mathcal{L}_{\text{p}}+\mathcal{L}_{\text{d}}+\mathcal{L}_{\text{s}}$ 13.767 0.08 0.06 0.06 9.153 2.87 0.92 0.47
IV-D Ablation Studies
To evaluate the effectiveness of the individual loss components introduced in Section III-C, we conduct a series of ablation experiments using different combinations of loss functions on our synthetic benchmark scenes. Note that the photometric loss serves as the foundation of our optimization framework. Without it, the network fails to converge to a meaningful solution.
As shown in Table IV, while the photometric loss alone enables the recovery of coarse geometric structures, it tends to overlook fine-scale surface variations, resulting in over-smoothed depth maps and missing details, especially near sharp edges and object boundaries. The inclusion of the surface color loss addresses this issue by encouraging the consistency between the predicted surface point color and the observed color, thus reinforcing the network’s ability to model subtle geometric details more faithfully.
Moreover, the distortion loss, which promotes a unimodal volume density distribution along each ray, helps to eliminate ambiguous or noisy depth estimates caused by multiple semi-transparent surfaces or low-frequency variations. By enforcing a single dominant depth response per ray, it significantly enhances the accuracy and stability of the final depth map. Together, the three loss terms work synergistically to improve both global structure recovery and local geometric fidelity.
Figure 10: Qualitative results under different numbers of projected patterns. The number below each image indicates how many patterns were used during both scene capture and voxel grid training. The rightmost image shows the ground truth depth map for reference.
We also examine the effect of the number of projected patterns on the performance of our method. In the full configuration, we generate a set of nine 2D binary patterns using unit square lengths of 20, 10, and 5 pixels. To evaluate the influence of pattern count, we progressively reduce the number of projected patterns by following a structured removal order: one pattern of length 20, followed by one of length 10, then one of length 5, and repeating this cycle (i.e., removal sequence: 20, 10, 5, 20, 10, 5, 20). This strategy ensures a balanced degradation of spatial information across different scales, allowing us to assess how various frequency components contribute to the depth estimation process. We verify 10 synthetic scenes for this experiment and summarize the results in Table V and Fig. 10. The results demonstrate that our voxel-based training process achieves satisfactory performance with as few as 6 patterns, with no significant improvement observed by further increasing the number of patterns. Notably, even with only four projected patterns, our approach still delivers promising results, highlighting its robustness and efficiency under limited input conditions.
TABLE V: Ablation Study on the Number of Projected Patterns in Our Framework
Pattern Count 2 3 4 5 6 7 8
MAE(mm) 203.9 31.7 18.3 16.9 15.3 15.2 15.0
V Conclusion